OpenStudy (anonymous):

A metal stays solid at Fahrenheit temps below 1778.3 degrees. Determine (in terms of an inequality) those Celsius temperatures for which the metal stays solid. Use the formula F=9/5 C +32. The solution set is {C|C < ? degrees}

OpenStudy (anonymous):

Hint: Rearrange the Farenheit formula. Plug in 1778.3 for F. Now write the inequality.

OpenStudy (anonymous):

I just don't get it...

OpenStudy (anonymous):

We want the "C" in the farenheit formula to be by itself. So the first thing you might want to do is subtract 32 from both sides of the equation.

OpenStudy (anonymous):

So then we would have 1746.3 = 9/5??

OpenStudy (anonymous):

Good try. When you rearranged the formula did you do it this way....... F = 9/5 C + 32 -32 -32 F - 32 = 9/5 C Multiply both sides by the reciprocal of 9/5(which is 5/9) 5/9(F - 32) = C so now we have the C by itself.

OpenStudy (anonymous):

I subtracted 32 from both sides and that's as far as I got. Fractions are very confusing for me.

OpenStudy (anonymous):

You are not alone my dear! But they are the foundation for algebra. If you have time a review is well advised. Anyway on with this problem:- Plug in 1746.3 in for the F. C = 5/9(1746.3 - 32) = 952.39 to 2 decimal places.

OpenStudy (anonymous):

So, have you any idea how to use this number in the clue they gave us in the question? I always find a good guess is as good as any other idea in these situations. Half the time I end up right! :-))

OpenStudy (anonymous):

This is all so confusing for me. I don't understand how you guys do this.

OpenStudy (anonymous):

So the temperature would be {C|C < 952.39}? Or 952.4 since they want it rounded to the nearest tenth.

OpenStudy (anonymous):

Practice, practice, practice. I understand if you are having trouble with fractions though, it really does make math more difficult. There are great booklets in places like Wal-Nart that run through the basics of fractions in a fun way. May I suggest you try one of those?

OpenStudy (anonymous):

Oh excellent! You got it!

OpenStudy (anonymous):

Thank you, I know I need more practice. So I got it right? Whoo Hoo!!

OpenStudy (anonymous):

Well crap, it says that is not the right answer.

OpenStudy (anonymous):

Calculator error(human error more like it)LOL. That's why it's good to write down your steps. See up there where I put the wrong value in? Where I got 1746.3 is beyond me!

OpenStudy (anonymous):

Now you try it with the correct number in it for F.

OpenStudy (anonymous):

I got 1746.3 as well.

OpenStudy (anonymous):

OK, I did it again but with the 1778.3 for the F C = 5/9(1178.3 - 32) = 970.2. so C < 970.2 Is that what they say is correct?

OpenStudy (anonymous):

I am sure it would help if I had an agebraic calculator!

OpenStudy (anonymous):

LOL.

OpenStudy (anonymous):

Yes, that is correct. Thank you. I wish I could do this stuff without help, but I truly appreciate your time. Thank you again.

OpenStudy (anonymous):

you're very welcome.